# P.O. Box 15500, 00076 Aalto, Finland

### v1

@@ 6.3 Network-Level Analysis of the Influence Spreading Model in Random Networks
$\text{ER}(1000,0.01)$ 0.93 0.94 0.80 0.63 0.45 0.30 0.18 0.11 0.06
$\text{WS}(1000,10,0.5)$ 0.89 0.89 0.79 0.66 0.53 0.38 0.25 0.16 0.10
$\text{BA}(1000,5)$ 0.96 0.88 0.86 0.85 0.84 0.82 0.73 0.42 0.17
In Table 1, we see that the $\text{BA}(1000,5)$ network shows quite strong to moderate positive correlation values for higher edge probabilities, whereas the correlation has significantly weakened for the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks. Eventually, positive correlation becomes weak for the $\text{BA}(1000,5)$ network as well.
Interestingly, the extension of the Barabási-Albert model proposed by Holme and Kim produced a random network that maintained a strong correlation between ISM betweenness and standard betweenness even for the highest edge probability value $0.9$. Stronger positive correlation between ISM betweenness and standard betweenness for the edge probabilities $0.7,0.8$, and $0.9$ was the only notable difference of the Holme-Kim network compared to the $\text{BA}(1000,5)$ network in our experiments.
@@
In Table 2, the Pearson correlation coefficient between out-centrality and degree centrality is shown for all edge probabilities. For the edge probability value $0.1$, the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display a quite strong positive correlation, which then becomes weaker as the edge probability increases. For the $\text{BA}(1000,5)$ network, there is also a positive correlation. However, the non-linear relationship makes the correlation weaker.
@@
$\text{ER}(1000,0.01)$ 0.95 0.93 0.82 0.69 0.55 0.42 0.31 0.22 0.16
$\text{WS}(1000,10,0.5)$ 0.93 0.96 0.90 0.82 0.70 0.57 0.43 0.31 0.23
$\text{BA}(1000,5)$ 0.87 0.68 0.57 0.49 0.41 0.29 0.15 0.06 0.03
In Figure 14, out-centrality is compared to the standard closeness centrality for $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ random networks. Again, it is important to notice the changing scale of out-centrality values as the edge probability increases. Out-centrality and standard closeness centrality display a similar relationship to what was seen between out-centrality and degree centrality.
In Table 3, we see a positive correlation for low edge probabilities, which becomes weaker as the edge probability increases, and eventually the correlation is close to zero. As edge probability increases, $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks maintain a stronger positive correlation than $\text{BA}(1000,5)$ network. We see that for low edge probabilities, closeness centrality is a good predictor for out-centrality values. We can see that for low edge probabilities, the close and immediate neighborhood of a node determines its ability to influence other nodes (out-centrality), since the low influence probability decays significantly over longer walk distances.
@@
$\text{ER}(1000,0.01)$ 0.97 0.96 0.87 0.76 0.63 0.50 0.39 0.29 0.21
$\text{WS}(1000,10,0.5)$ 0.89 0.87 0.82 0.74 0.65 0.54 0.42 0.32 0.25
$\text{BA}(1000,5)$ 0.92 0.76 0.65 0.56 0.47 0.33 0.18 0.07 0.03
@@
In Figure 15, the relationship between in-centrality and out-centrality is studied for the $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ networks. It is important to note that both the x-axis and y-axis scale changes in every individual plot. We see the same behavior for out-centrality and in-centrality values that was seen for the Krackhardt kite and small network examples in Sections 6.1 and 6.2, respectively. Out-centrality values have much larger dispersion, whereas there is a small variation between in-centrality values.
In Figure 15 and in Table 4, both $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display an initially strong positive correlation, which eventually reverses into a strong negative correlation as the edge probability increases. However, for the $\text{BA}(1000,5)$ network, the reversal has already begun for the edge probability value $0.1$ since part of the scatter plot shows a positive correlation, while the other part shows a negative correlation. As the edge probability increases, the correlation between in-centrality and out-centrality becomes strongly negative. We are again essentially seeing the inverting ranking behavior of out-centrality and in-centrality as the edge probability increases.
@@
$\text{ER}(1000,0.01)$ 0.99 -0.98 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
$\text{WS}(1000,10,0.5)$ 0.99 -0.99 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
$\text{BA}(1000,5)$ 0.53 -0.97 -0.99 -1.00 -1.00 -1.00 -1.00 -1.00 -1.00
6.4 Special Cases
In Figure 16, a tetragon network is shown on the left and the same network with one peripheral node added is shown on the right. For the regular tetragon network, given an identical edge probability, each node attains the same ISM betweenness centrality value, the same out-centrality value, and the same in-centrality value. When a peripheral node is added, the results change in a way that node 4 becomes the highest ranked in ISM betweenness value, whereas node 5 behaves like the peripheral nodes we saw in our earlier two examples. Nodes 1 and 3 still attain identical values to each other, although their numerical values change.
4132
41325
Figure 16: A tetragon network on the left. When a peripheral node is added on the right, the results of the influence spreading model change dramatically.
The tetragon network shows that not all networks produce the inverting ranking behavior between out-centrality and in-centrality as was seen for the Krackhardt kite network in Figure 4, for the small node network in Figure 8 and for the larger random networks. It also further reinforces that when the network contains symmetry, the results of the influence spreading model are also symmetric, as long as the edge probabilities are the same for all edges.
Another case, when network-wide inverting ranking behavior between out-centrality and in-centrality does not happen, is when nodes are completely disconnected from the rest of the network. If a node is disconnected from the rest of the network i.e. it has degree zero, then its in-centrality and out-centrality equal zero for all edge probabilities. The rest of the network may exhibit a reversing behavior between in-centrality and out-centrality; however, nodes without neighbors disrupt the change in correlation from positive to negative, as their values remain static at zero.
A particular Dutch student network is an example of this phenomenon, where the correlation between in-centrality and out-centrality remains positive as the edge probability increases. The Dutch student network consists of 32 students, where their friendship was originally mapped and studied in [van1999], and the network is visualized in Figure 17. This student network contains three students (nodes) whose degree is zero. In actuality, the network has four components: one very large component and the three separated nodes, each forming its own component.
2621147162102231817242331525321911276192829301320451218
Figure 17: A Dutch student network with three nodes (5, 12, and 18) completely disconnected from all other nodes. Based on data collected by Van de Bunt [van1999]
In Figure 18, we see that although the rest of the nodes in the network display the reversing behavior between in-centrality and out-centrality as the edge probability increases, the three zero-degree nodes maintain their position as last in the rankings. If the nodes with degree zero were removed, we would observe the same inverted ranking behavior for the whole network that was seen before.
Figure 18: Out-centrality and in-centrality rankings for the Dutch student network shown in Figure 17. Node labels are plotted alongside the lines. Notably, multiple nodes share the same rank.
In Figure 19, the in-centrality and out-centrality values for this student network are plotted in a scatter plot. We can see that the positive correlation persists because there are three zero-degree nodes with both in-centrality and out-centrality values of zero.
Figure 19: In-centrality and out-centrality values plotted for increasing edge probabilities for the Dutch student network shown in Figure 17
In Table 5, we observe a positive correlation between in-centrality and out-centrality values across all edge probabilities. For edge probabilities in the middle range, the correlation becomes weaker but remains positive. However, for the highest edge probability, the correlation strengthens significantly, becoming strongly positive again. If we remove the three zero-degree outlier nodes, the correlation shifts from positive to negative, as illustrated by the ranking changes in Figure 19.
@@ 6.4 Special Cases
Students 1.00 0.98 0.63 0.48 0.60 0.75 0.87 0.95 0.99
Table 5: Pearson correlation coefficient between in-centrality and out-centrality for the Dutch student network shown in Figure 17.
@@ 7 Discussion
The influence spreading model serves as a versatile framework to investigate properties of various complex networks as it also allows for directed unidirectional or bidirectional edges and any edge probabilities to be used. In this article, however, we have studied undirected networks, where all edge probabilities are assumed to be equal. This setting has given us a controlled environment, where the effect of the edge probability can be captured more easily. We have also developed an understanding of how the influence spreading model behaves when the edge probability is varied.
@@
All experiments performed in Sections 6.1, 6.2, and 6.3 were performed using the maximum allowed walk length $L_{max}=20$. We also performed all these experiments for shorter and longer maximum allowed walk lengths. The effect of the maximum length of allowed walk is that the reversing ranking behavior in the in-centrality occurs earlier if walks are allowed to be longer than the length of $20$, and reversing occurs later if walks are only allowed to be shorter than the length of $20$. In general, a longer $L_{max}$ increases the probability of influencing others, as it allows more walks within the network. In [kuikka2022efficiency], the effect of varying the maximum allowed walk length was discussed in detail (the article used the term path instead of walk while retaining the meaning of walk as defined in this article).
@@
In this study, we have only presented the complex contagion version of the influence spreading model. For random networks produced by the Erdős-Rényi, Watts-Strogatz, and Barabási-Albert models, the complex contagion process results in in-centrality and out-centrality values being positively correlated for low edge probabilities. As the edge probability is increased, eventually this relationship between in-centrality and out-centrality completely reverses, where there exists a negative correlation between these two measures. It is important to emphasize that simple contagion cannot create this phenomenon since cycles are not allowed, and therefore there are the same number of paths from a source node to a target node as there are from the target node to the source node. In complex contagion processes, cycles act as an amplifier. In social networks, this amplification effect could be interpreted as an echo chamber for ideas.
We have also seen how in-centrality and out-centrality characterize the nodes of the network when everything else is assumed to be equal, i.e. when the edges of the network are undirected and when the edge probabilities are the same for all edges of the network. In this case, the network topology determines the in-centrality and out-centrality values of the nodes. We have seen how peripheral nodes tend to be the least influential nodes (lower out-centrality values), whereas more central nodes have more influence (higher out-centrality values). This phenomenon seems to persist as edge weight increases. We have also seen how, for low edge probabilities, peripheral nodes are influenced the least due to lower in-centrality values, whereas more central nodes are being influenced the most because of higher in-centrality values. However, as the edge weight increases, at some point the situation reverses itself, and peripheral nodes become the most influenced nodes, and the most central nodes become the least influenced nodes. In real social networks, for example, it is plausible to think that the peripheral nodes are the least influential nodes, whereas the most central nodes are the most influential nodes. Additionally, in real social networks, nodes can be so isolated that they are influenced very little by the rest of the network due to their isolation, whereas more central nodes are influenced by at least their immediate neighborhood. The in-centrality and out-centrality defined by the influence spreading model produce behavior that aligns well with the aforementioned behavior of a real social network.
The proposed ISM betweenness centrality has been shown to be capable of revealing what standard betweenness centrality cannot. For example, the ISM betweenness can highlight the importance of alternative routes in a network. It also emphasizes nodes that act as important bridges in a network. However, the ISM betweenness centrality has a great computational cost. Calculating the ISM betweenness centrality requires that the influence spreading algorithm be run again for every single node, since removing a different node from the network creates a new network. The ISM betweenness centrality reveals more of the network topology than the standard betweenness centrality, but this advantage comes at a greater computational cost.
@@ 8 Conclusion
Many standard centrality measures, such as closeness and betweenness centrality, are based on shortest paths in the network. The influence spreading model defines centrality metrics based on unrestricted paths, walks, and cycles. We have seen how there are similarities between the centrality metrics defined by the influence spreading model and standard centrality metrics, especially for low edge probabilities. We have also seen how the centrality metrics defined by the influence spreading model can reveal information that centrality metrics based on shortest paths cannot. The novelty presented in this article is the detailed examination of the behavior of ISM betweenness centrality, out-centrality, and in-centrality in comparison to standard centrality metrics.
We have demonstrated that ISM betweenness centrality can reveal the importance of alternative routes while maintaining similarity to standard betweenness centrality. We have also demonstrated that out-centrality can detect central nodes as influential nodes and in-centrality can detect peripheral nodes as the most influenced nodes when all else is equal (network edges are undirected and all edge probabilities are equal). These results give validity to out-centrality and in-centrality measures, since we achieve results that are logical. Therefore, we can infer that when studied networks are made more realistic with directed edges and different edge probabilities, we will see even clearer differences between the roles of nodes when our influence-based centrality metrics are used.

### v2

@@ 4.2 Network-Level Analysis of the Influence Spreading Model in Random Networks
In Table 1 we see that the $\text{BA}(1000,5)$ network shows quite strong to moderate positive Pearson correlation values for higher edge probabilities, whereas for the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks the correlation has significantly weakened. For the $\text{BA}(1000,5)$ network and for the highest edge probabilities, Pearson correlation becomes weak as well.
@@
$\text{ER}(1000,0.01)$ 0.929 0.939 0.803 0.630 0.451 0.296 0.182 0.107 0.061
$\text{WS}(1000,10,0.5)$ 0.887 0.891 0.787 0.665 0.527 0.380 0.250 0.157 0.105
$\text{BA}(1000,5)$ 0.964 0.883 0.857 0.847 0.839 0.819 0.727 0.424 0.171
In Table 2, we see that for the $\text{BA}(1000,5)$ network Spearman’s rank correlation remains quite high for all edge probabilities. However, for the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks Spearman’s rank correlation becomes weaker as edge probability increases. A notable difference to Pearson correlation is that for the highest edge probabilities there is more correlation between the ranks than there is between the values.
Table 2: Spearman’s rank correlation coefficient between ISM betweenness centrality and standard betweenness centrality.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.936 0.979 0.877 0.721 0.590 0.506 0.456 0.437 0.429
$\text{WS}(1000,10,0.5)$ 0.871 0.887 0.796 0.678 0.553 0.446 0.363 0.313 0.269
$\text{BA}(1000,5)$ 0.915 0.923 0.911 0.896 0.878 0.860 0.844 0.834 0.831
Interestingly, the extension of the Barabási-Albert model proposed by Holme and Kim produced a random network that maintained a strong Pearson correlation between ISM betweenness and standard betweenness even for the highest edge probability value $0.9$. Stronger positive Pearson correlation between ISM betweenness and standard betweenness for the edge probabilities $0.7,0.8$, and $0.9$ was the only notable difference of the Holme-Kim network compared to the $\text{BA}(1000,5)$ network in our experiments.
@@
In Table 3, the Pearson correlation coefficient between out-centrality and degree centrality is shown for all edge probabilities. For the edge probability value $0.1$, the $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display a quite strong positive Pearson correlation, which then becomes weaker as the edge probability increases. For the $\text{BA}(1000,5)$ network, there is also a positive Pearson correlation. However, the non-linear relationship makes the Pearson correlation weaker. In Table 4, we see that Spearman’s rank correlation remains high for all random networks as the edge probability increases. We can infer that as edge probability increases for all edges simultaneously, the relative rankings between nodes do not change much.
@@
$\text{ER}(1000,0.01)$ 0.950 0.928 0.821 0.690 0.552 0.422 0.310 0.222 0.157
$\text{WS}(1000,10,0.5)$ 0.935 0.962 0.905 0.817 0.702 0.567 0.431 0.315 0.232
$\text{BA}(1000,5)$ 0.865 0.681 0.573 0.491 0.406 0.289 0.154 0.063 0.030
Table 4: Spearman’s rank correlation coefficient between out-centrality and degree centrality.
@@
$\text{ER}(1000,0.01)$ 0.947 0.992 0.995 0.995 0.995 0.995 0.995 0.995 0.997
$\text{WS}(1000,10,0.5)$ 0.928 0.987 0.987 0.987 0.987 0.987 0.987 0.987 0.999
$\text{BA}(1000,5)$ 0.848 0.976 0.985 0.985 0.985 0.985 0.985 0.985 0.984
In Figure 10, the out-centrality is compared to the standard closeness centrality for $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ random networks. Again, it is important to notice the changing scale of out-centrality values as the edge probability increases. Out-centrality and standard closeness centrality display a similar relationship as seen between out-centrality and degree centrality.
In Table 5, we see a positive Pearson correlation for low edge probabilities, which becomes weaker as the edge probability increases, and eventually the correlation is close to zero. As edge probability increases, $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks maintain a stronger positive Pearson correlation than $\text{BA}(1000,5)$ network. We see that for low edge probabilities, closeness centrality is a good predictor for out-centrality values. We can see that for low edge probabilities, the close and immediate neighborhood of a node determines its ability to influence other nodes (out-centrality), since the low influence probability decays significantly over longer walk distances.
@@
$\text{ER}(1000,0.01)$ 0.974 0.959 0.871 0.757 0.630 0.502 0.385 0.288 0.212
$\text{WS}(1000,10,0.5)$ 0.893 0.873 0.817 0.745 0.652 0.541 0.425 0.323 0.248
$\text{BA}(1000,5)$ 0.917 0.762 0.649 0.561 0.467 0.335 0.179 0.073 0.033
In Table 6, we can see that Spearman’s rank correlation remains relatively high compared to Pearson correlation for all random networks as the edge probability increases. Interestingly, when edge probabilities are high, the $\text{BA}(1000,5)$ network has lower rank correlation than the other two networks.
In Figure 11, the relationship between in-centrality and out-centrality is studied for the $\text{ER}(1000,0.01)$, $\text{WS}(1000,10,0.5)$, and $\text{BA}(1000,5)$ networks. It is important to note that both the x-axis and y-axis scale changes in every individual plot. We see the same behavior for out-centrality and in-centrality values that was seen for the Krackhardt kite network in Section 4.1. Out-centrality values have much larger dispersion, whereas there is a small variation between in-centrality values.
Table 6: Spearman’s rank correlation coefficient between out-centrality and closeness centrality.
@@
$\text{ER}(1000,0.01)$ 0.991 0.976 0.968 0.965 0.963 0.962 0.961 0.960 0.952
$\text{WS}(1000,10,0.5)$ 0.892 0.881 0.877 0.874 0.872 0.869 0.867 0.866 0.849
$\text{BA}(1000,5)$ 0.901 0.735 0.701 0.693 0.687 0.684 0.682 0.682 0.673
In Figure 11 and in Table 7, both $\text{ER}(1000,0.01)$ and $\text{WS}(1000,10,0.5)$ networks display an initially strong positive Pearson correlation, which eventually reverses into a strong negative Pearson correlation as the edge probability increases. However, for the $\text{BA}(1000,5)$ network, the reversal has already begun for the edge probability value $0.1$ since part of the scatter plot shows a positive correlation, while the other part shows a negative correlation. As the edge probability increases, the correlation between in-centrality and out-centrality becomes strongly negative. We are again essentially seeing the inverting ranking behavior of out-centrality and in-centrality as the edge probability increases. In Table 8, this inverted ranking can be verified by looking at the change in rank correlation as edge probability increases.
In our experiments we have discovered that network-wide inverting ranking behavior between out-centrality and in- centrality does not occur when some nodes are completely disconnected from the rest of the network. For such a disconnected (zero degree) node, its in-centrality and out-centrality equal zero for all edge probabilities. The rest of the network may exhibit a reversing behavior between in-centrality and out-centrality, although nodes without neighbors disrupt the change in correlation from positive to negative as their values remain static at zero.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 0.995 -0.984 -0.998 -0.999 -1.000 -1.000 -1.000 -1.000 -1.000
$\text{WS}(1000,10,0.5)$ 0.992 -0.987 -0.998 -1.000 -1.000 -1.000 -1.000 -1.000 -1.000
$\text{BA}(1000,5)$ 0.532 -0.972 -0.995 -0.999 -1.000 -1.000 -1.000 -1.000 -1.000
Table 8: Spearman’s rank correlation coefficient between in-centrality and out-centrality.
Edge probability
Network 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$\text{ER}(1000,0.01)$ 1.000 -0.989 -0.999 -1.000 -1.000 -1.000 -1.000 -0.999 -0.966
$\text{WS}(1000,10,0.5)$ 0.994 -0.992 -1.000 -1.000 -1.000 -1.000 -1.000 -0.992 -0.976
$\text{BA}(1000,5)$ 0.828 -0.958 -0.999 -1.000 -1.000 -1.000 -1.000 -0.999 -0.990
@@ 5 Discussion
The influence spreading model provides a flexible framework for studying a wide range of complex networks, as it supports directed and undirected edges as well as arbitrary edge probabilities. In this study, we focus on undirected networks with equal edge probabilities, which provides a controlled setting to isolate the effects of edge probability variation and understand the behavior of the model as this parameter changes.
@@
All experiments in Sections 4.1 and 4.2 were conducted using a maximum walk length of $L_{max}=20$. Additional experiments with shorter and longer walk lengths show that allowing longer walks causes the reversing behavior in in-centrality rankings to occur earlier, whereas restricting walk length delays this effect. In general, increasing $L_{max}$ raises the probability of influence by permitting a greater number of walks. The impact of varying the maximum allowed walk length has been discussed in detail in [21], where the term path was used with the same meaning as walk in this article.
@@
This study considers only the complex contagion version of the influence spreading model. For random networks generated using the Erdős–Rényi, Watts– Strogatz, and Barabási–Albert models, in-centrality and out-centrality are positively correlated at low edge probabilities. As the edge probability increases, this relationship reverses and becomes negative. This phenomenon cannot occur under simple contagion, where cycles are absent and the number of paths between node pairs is symmetric. In complex contagion, cycles act as amplifiers, which in social networks can be interpreted as echo chambers.
When networks are undirected and edge probabilities are identical, the in-centrality and out-centrality are determined solely by network topology. Peripheral nodes tend to have lower out-centrality, while more central nodes exhibit greater influence, a pattern that persists as edge probability increases. For low edge probabilities, the peripheral nodes are also the least influenced, whereas the central nodes are the most influenced. As edge probability increases, this pattern reverses, with peripheral nodes becoming the most influenced and central nodes the least influenced. This behavior is consistent with the intuitive properties of real social networks, where the central nodes are typically more influential and more exposed to influence.
The proposed ISM betweenness centrality reveals structural properties that standard betweenness centrality cannot reveal, such as the importance of alternative routes and critical bridging nodes. However, this increased expressiveness comes at a high computational cost, as the influence spreading algorithm must be recomputed for each node removal. Thus, while ISM betweenness captures more of the underlying network topology, it does so at the expense of computational efficiency.
Our study reveals similar behavior between the centrality measures proposed by the influence spreading model and the standard centrality measures. However, the influence spreading model also allows for in-built probability interpretation for the values of centrality measures. In some application areas, this natural probability interpretation can be more valuable than the values provided by other standard centrality measures. Our study also highlights that the values and rankings calculated from different centrality measures would reveal different information.
@@ 6 Conclusion
Many standard centrality measures, including closeness and betweenness centrality, rely on shortest paths. In contrast, the influence spreading model defines centrality metrics based on unrestricted paths, walks, and cycles. For low edge probabilities, the resulting metrics exhibit similarities to standard centrality measures, while also revealing structural information that shortest-path-based metrics cannot capture. The main contribution of this study is a detailed analysis of ISM betweenness centrality, out-centrality, and in-centrality compared to standard centrality measures.
We showed that ISM betweenness centrality highlights the importance of alternative routes while remaining comparable to standard betweenness centrality. In addition, out-centrality identifies influential central nodes, whereas in-centrality highlights peripheral nodes as the most influenced when network edges are undirected and edge probabilities are uniform. These results support the validity of the proposed centrality measures and suggest that in more realistic networks with directed edges and heterogeneous edge probabilities, the distinctions between node roles will become even more pronounced.